OpenStudy (anonymous):

Find the slope of the line that passes through the pair of points. (1, 7), (10, 1)

OpenStudy (studygurl14):

Use the slope formula. Do you know what that is? \(\large\tt And...\) \(\Large\sf\color{blue}{Welcome~To~OpenStudy!}\)

Nnesha (nnesha):

@roro500 ???

OpenStudy (anonymous):

X1 y1. X2 y2 (1, 7). (10, 1) Y2-y1 --------- X2-x1 M= 1-7 ----- =-6/9 = -0.6 10-1.